# SwiftKV: Fast Prefill-Optimized Inference with Knowledge-Preserving Model Transformation

### v1

@@ 5.3 The impact of fine-tuning datasets
The results are shown in Table 4 (a). The original Llama-3.1-8B-Instruct has a average score of 73.71 but the model trained using our two datasets only achieved 65.77. This indicates the training data used for SwiftKV is not optimal and there may be opportunities to further improve the results we reported in Sec. 4 as discussed next.
@@
The results are reported in Table 4 (b). The performance of all tasks except Winogrande are slightly improved, with the average score being 0.23 higher. Particularly, GSM-8K improves the most, with a 0.53% improvement. This is expected since we added extra math and coding datasets. Considering the small amount of new data (83k vs. 1.2M), the improvement is remarkable.
@@
Table 4: The impact of datasets on Llama-3.1-8B-Instruct.
@@
5.4 Combining with other KV cache compression methods
SwiftKV explores an orthogonal design space from many existing KV cache compression methods, which means that it can be easily combined with them, e.g., sliding window (Jiang et al., 2023), token-level pruning (Liu et al., 2024d), quantization (Hooper et al., 2024) etc. In this section, we show the combined effect of SwiftKV with per-token KV cache FP8 quantization  (Yao et al., 2022) using PyTorch’s natively supported float8_e4m3fn.  Table 5 shows the accuracy degradation is within 0.4 points for all cases, even though we applied post-training quantization with no quantization-aware finetuning.
Appendix D explores a second, potentially interesting, trade-off between AcrossKV (inter-layer) vs GQA (intra-layer) KV cache sharing.
Table 5: Llama-3.1-8B-Instruct KV cache quantization results.
Figure 5: Density of early exit probabilities and alignment of early exit vs final logits.
@@ 3. Yellow
Table 6: A Q&A example of early exit.
5.5 Simple Early Exit for Decoding Tokens
@@
To test the feasibility, we added an early exit language modeling head. We then used the input to SingleInputKV layer to calculate output logits, and incorporated them as part of the distillation training. Our results are preliminary and requires further evaluation, but we found that the alignment between early exit logits and the final output logits to be over 66% when the largest probability from the early exit logits is over 95% (Fig. 5). We used this as our early exit criteria (i.e., decoding tokens exit early if it predicts an output token with 95%+ probability), and  Table 6 shows a sample result. See Appendix E for more details.
6 Conclusions
In this paper, we presented SwiftKV, a novel model transformation for reducing inference cost for prompt-dominant workloads, combined with a KV cache reduction strategy to reduce memory footprint, and a light-weight distillation procedure to preserve model accuracy. While we presented strong results on the effectiveness of SwiftKV, exploration of parameter-preserving model transformations for inference optimization is still in its early stages. We have identified both limitations as well as areas of improvement. Given the simplicity and effectiveness of SwiftKV, we hope that this will spark further exploration from the AI community which we hope will continue lower the latency and cost of inference.
Limitations and Future Work
It is important for every work to acknowledge its limitations and suggest future directions, particularly for LLM-related works. In our work, we did not aim to optimize the training data selection though we provide potential ways in Sec. 5.3. Additionally, we did not include a detailed benchmark analysis for our method. However, as shown in Sec. 5.3, we ensured that our datasets were not cherry-picked to overfit the reported tasks. Furthermore, we did not finetune our model with advanced post-training approaches, like DPO and RLHF, which we leave for future work. Finally, we hypothesize that our method can work even better when combined with pretraining or continued-pretraining, but due to resources constraints, we did not explore this direction. We hope to revisit these ideas in the future.
Appendix A Additional Motivation
Appendix B Calculating Computation Reduction
For a vanilla transformer model Vaswani et al. (2017) with a hidden dimension size $d$, the total amount of computation per token per layer is roughly given by $24d^{2}$, where $4d^{2}$ goes to KV projections, another $4d^{2}$ goes to Q projection and context GEMM, and the remaining $16d^{2}$ goes to MLP computation. For popular GQA (Ainslie et al., 2023b) based architectures like Llama (Touvron et al., 2023), Mistral (Jiang et al., 2023), Qwen (Bai et al., 2023), and many others, the KV projection is reduced to $4d^{2}/G$, where $G$ is the group size and is $\geq 4$. As such, the compute cost of the KV projection is negligible ($\frac{4d^{2}/G}{20d^{2}+4d^{2}/G}$ which is less than 5% when $G\geq 4$), compared to the rest of the computation.
With SingleInputKV, pre-fill tokens only require computing this negligible KV projection for layers $>l$. Therefore, it reduces the total pre-fill computation to approximately $l/L$. Fig. 2 shows significant end-to-end reduction in model forward time for different value of $l$ compared to $L$. In Sec. 4.3, we show that such compute reduction translates to considerably higher aggregate serving throughput and lower latency, reducing both serving cost, and improving end-user experience.
Appendix C Experimental Details
C.1 Training and evaluation details
For all tasks, we follow the same number of few shots and/or chain of thoughts as the provided commands. We present the number of shots and metrics used in the paper in Table C.1.
Table C.1: The setting for different tasks
C.2 Inference Speedup Evaluation Details
Figure C.1: Time to first token (TTFT, top) and time per output token (TPOT, bottom) for input lengths 2000 (left), 8000 (middle), and 32000 (right) for Llama-3.1-8B. For each experiment, a range of different request arrival rates is simulated. Each request generates 256 output tokens.
Hardware Details.
vLLM Configuration.
Llama-3.1-8B Latency Evaluation.
See Fig. C.1.
Appendix D Inter-layer AcrossKV vs Intra-Layer KV cache Reduction
Table D.1: Llama-3.1-8B-Instruct AcrossKV design
We present the $50\%$ SingleInputKV reduction with MQA, GQA plus AcrossKV, and GQA plus MHA in Table D.1, that all have the same KV cache reduction, 37.5%. AcrossKV-GQA actually provides the best performance. One thing to notice is that the AcrossKV-MHA is actually worse than the result of AcrossKV-16x from from Table 1 even though AcrossKV-MHA has larger KV cache than AcrossKV-16x. We hypothesis that this might be related to hyper-paramter tuning but did not invest deeper. Also, note that pure MQA leads to worst performance, which is about 17 points lower than AcrossKV-GQA
Appendix E Early Exit Details
@@
After training, we first look the alignment between the early exit logits and the final logits. The figure is shown in Fig. 5. The left axis plot the density of early exit largest probabilities. And the right axis shows the probabilities when the maximum of early exit logits aligns with the final logits for each bin. As can be seen, when the maximum logits is larger than $0.95$, the alignment likelihood is larger than 66%. Note that this is achieved within 160M token training and longer training should help the alignment.
How to use early exit is always an interesting directly and research topic. Fully exploring this is out of the scope of this paper. Here we adopt a simple heuristic approach. When the largest probability is larger than 0.95, we directly use the early exit prediction. Otherwise, we still process the remaining layers to get the final prediction. We apply this logic to some questions we selected from Alpaca (Taori et al., 2023) and the results are shown in Table 6 and Appendix E.1.
E.1 Early Exit Examples

### v2

@@ Appendix D The impact of fine-tuning datasets
The results are shown in Table D.1 (a). The original Llama-3.1-8B-Instruct has a average score of 73.71 but the model trained using our two datasets only achieved 65.77. This indicates the training data used for SwiftKV is not optimal and there may be opportunities to further improve the results we reported in Sec. 4 as discussed next.
@@
The results are reported in Table D.1 (b). The performance of all tasks except Winogrande are slightly improved, with the average score being 0.23 higher. Particularly, GSM-8K improves the most, with a 0.53% improvement. This is expected since we added extra math and coding datasets. Considering the small amount of new data (83k vs. 1.2M), the improvement is remarkable.
@@
Table D.1: The impact of datasets on Llama-3.1-8B-Instruct.
@@
Figure D.1: Density of early exit probabilities and alignment of early exit vs final logits.
@@ 3. Yellow
Table D.2: A Q&A example of early exit.
Appendix E Simple Early Exit for Decoding Tokens
@@
To test the feasibility, we added an early exit language modeling head. We then used the input to SingleInputKV layer to calculate output logits, and incorporated them as part of the distillation training. Our results are preliminary and requires further evaluation, but we found that the alignment between early exit logits and the final output logits to be over 66% when the largest probability from the early exit logits is over 95% (Fig. D.1). We used this as our early exit criteria (i.e., decoding tokens exit early if it predicts an output token with 95%+ probability), and  Table D.2 shows a sample result. See Appendix F for more details.
Appendix F Early Exit Details
@@
After training, we first look the alignment between the early exit logits and the final logits. The figure is shown in Fig. D.1. The left axis plot the density of early exit largest probabilities. And the right axis shows the probabilities when the maximum of early exit logits aligns with the final logits for each bin. As can be seen, when the maximum logits is larger than $0.95$, the alignment likelihood is larger than 66%. Note that this is achieved within 160M token training and longer training should help the alignment.
How to use early exit is always an interesting directly and research topic. Fully exploring this is out of the scope of this paper. Here we adopt a simple heuristic approach. When the largest probability is larger than 0.95, we directly use the early exit prediction. Otherwise, we still process the remaining layers to get the final prediction. We apply this logic to some questions we selected from Alpaca (Taori et al., 2023) and the results are shown in Table D.2 and Appendix F.1.
F.1 Early Exit Examples